# Cognitive Hierarchy in Day-to-day Network Flow Dynamics

### v1

@@ Appendix C Derivation of Fact 1
Denote $\boldsymbol{z}^{k,(t)}$ as $\boldsymbol{x}^{k,(t)}-\gamma\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right),k=0,1,2$. For the 0-step travelers, take the Jacobian, $J[\cdot]$, for RHS of (9) w.r.t. $\boldsymbol{x}^{0,(t)}$:
@@
where $J\left[\frac{\boldsymbol{c}\left(\boldsymbol{\pi}^{1,(t+1)}\right)}{\boldsymbol{x}^{k,(t)}}\right],k=0,1,2,$ is given by (C).
@@
where $J\left[\frac{\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)}{\boldsymbol{x}^{k,(t)}}\right],k=0,1,2,$ is given by (C).
@@ Appendix D Proof of Proposition 3
According to Proposition 2, at the DUE, $\boldsymbol{\pi}^{0,(t+1)}=\boldsymbol{\pi}^{1,(t+1)}=\boldsymbol{\pi}^{2,(t+1)}$ and thus the Jacobians of route cost function $D_{2}=D_{1}=D_{0}=D_{\ast}$, where $D_{\ast}$ is the Jacobian of the route cost function evaluated at the DUE $\tilde{\boldsymbol{x}}^{\ast}$. As per the assumption that the projection operator always yields positive route flow, all the $Q$’s and $\hat{Q}$’s in the Jacobian (22) reduce to $\bar{Q}$ (defined in Lemma 2). Moreover, since $\alpha=\hat{\alpha}=1$ and $\hat{\gamma}=\gamma$, each sub-block matrix in (22) reduces to:
@@
Denote $A$ as $\bar{Q}(I-\gamma D_{\ast})$. By the identity that $\bar{Q}A=\bar{Q}\bar{Q}(I-\gamma D_{\ast})=A$ (see Lemma 2) and some matrix algebraic calculations, Jacobian (22) can be further represented as:
@@
Readers can verify that $A\cdot(A-\bar{Q})A=(A-\bar{Q})A\cdot A$ using the identity $\bar{Q}A=\bar{Q}$, i.e., $\lambda I-A$ commutes with $-(A-\bar{Q})A$. Referring to Theorem A.2, we have
@@
(D) shows that for $|K|=2$, $JP$ has $2\cdot\sum_{w}{|R_{w}|}$ eigenvalues, half of which are the eigenvalues of $\bar{Q}$ and the other half are the ones of $A^{2}$. For each OD pair $w$, we know from Lemma 2 that $\bar{Q}_{w}$ has $|R_{w}|$ eigenvalues with $|R_{w}|-1$ of them being 1 and the other one being 0. So $JP$ will have $\sum_{w}{(|R_{w}|-1)}$ eigenvalues being 1 and $|W|$ ones being 0. Due to the geometric and algebraic multiplicities with eigenvalue one both being $|R_{w}|-1$, every eigenvalue on the unit circle has an associated Jordan block of order 1, which satisfies the condition (ii) in Theorem A.1. As a result, the stability only depends on the eigenvalues of $A^{2}$.
@@
Row and column operations do not change the determinant. Subtracting the third “block column” of the determinant from the first and second block columns and then adding the first and second “block rows” to the third block row, the determinant in (D.5) becomes an upper triangular matrix:
@@ Appendix E Proof of Lemma 3
$Q\left[\boldsymbol{z}\right](I-\gamma D\left[\boldsymbol{x}\right])$ is not diagonalizable because $Q\left[\boldsymbol{z}\right](I-\gamma D\left[\boldsymbol{x}\right])$ does not commute with $-\gamma Q\left[\boldsymbol{z}\right]D\left[\boldsymbol{x}\right]$. Still, $Q\left[\boldsymbol{z}\right](I-\gamma D\left[\boldsymbol{x}\right])$ share the same eigenvector with $-\gamma Q\left[\boldsymbol{z}\right]D\left[\boldsymbol{x}\right]$ for nonzero eigenvalues. Suppose $\mu_{i}$ is a nonzero eigenvalue of $Q\left[\boldsymbol{z}\right]D\left[\boldsymbol{x}\right]$ with corresponding eigenvector $\boldsymbol{v}$, i.e., $Q\left[\boldsymbol{z}\right]D\left[\boldsymbol{x}\right]\boldsymbol{v}=\mu_{i}\boldsymbol{v},\mu_{i}\neq 0$. Then we have from (E) that $(-\gamma\xi_{i}\mu_{i})\boldsymbol{v}=(-\gamma\mu_{i}+\gamma^{2}\mu_{i}^{2})\boldsymbol{v}$ where $\xi_{i}$ is the corresponding eigenvalue for $Q\left[\boldsymbol{z}\right](I-\gamma D\left[\boldsymbol{x}\right])$. Solving it yields $\xi_{i}=1-\gamma\mu_{i},\mu_{i}\neq 0$.
@@ Appendix F Proof of Proposition 4
where $A=\bar{Q}(I-\gamma D_{\ast})$ and $\hat{A}=\bar{Q}(I-\hat{\gamma}D_{\ast})$. Referring to Theorem A.2, we have $|(\lambda I-A\hat{A}+\hat{A}-A)||(\lambda I-\bar{Q})|=0$. $A$ and $\hat{A}$ share the same eigenvector $\boldsymbol{v}$ with $\bar{Q}D_{\ast}$ for the eigenvalue $\lambda$ of $\bar{Q}D_{\ast}$ being nonzero. Assume $\bar{Q}D_{\ast}\boldsymbol{v}=\lambda\boldsymbol{v},\lambda\neq 0$. The matrix $A\hat{A}-\hat{A}+A$ determining the stability multiplied by $\boldsymbol{v}$ becomes:
@@
Appendix G Derivation of Fact 2
@@
For the 0-step travelers, take the Jacobian, $J[\cdot]$, of the dynamical system (25) w.r.t. $\boldsymbol{x}^{k,(t)}$:
@@
$=\alpha p^{0}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\alpha)I,$ (G.1)
@@
$=\alpha p^{0}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D[\tilde{\boldsymbol{x}}^{(t)}].$ (G.2)
@@
$=D\left[\boldsymbol{\pi}^{1,(t+1)}\right]\cdot\left(\hat{\alpha}\Upsilon^{\hat{\theta}}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\hat{\alpha})I\right).$ (G.3)
@@
$\quad=\alpha p^{1}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{1,(t+1)}\right)\right]\cdot D\left[\boldsymbol{\pi}^{1,(t+1)}\right]\cdot\left(\hat{\alpha}\Upsilon^{\hat{\theta}}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\hat{\alpha})I\right)+(1-\alpha)I,$ (G.4)
@@
$\quad=\alpha{p^{1}}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{1,(t+1)}\right)\right]\cdot D\left[\boldsymbol{\pi}^{1,(t+1)}\right]\cdot\left(\hat{\alpha}\Upsilon^{\hat{\theta}}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\hat{\alpha})I\right).$ (G.5)
@@
$\cdot\left(D\left[\boldsymbol{\pi}^{1,(t+1)}\right]\cdot\left(\hat{\alpha}\Upsilon^{\hat{\theta}}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\hat{\alpha})I\right)\right)+(1-\hat{\alpha})I.$ (G.6)
@@
$J\left[\frac{\alpha{p^{2}}\Phi^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)\right]+(1-\alpha)\boldsymbol{x}^{2,(t)}}{\boldsymbol{x}^{k\neq 2,(t)}}\right]=\alpha{p^{2}}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)\right]\cdot D\left[\boldsymbol{\pi}^{2,(t+1)}\right]\cdot J\left[\frac{\boldsymbol{\pi}^{2,(t+1)}}{\boldsymbol{x}^{k,(t)}}\right],$ (G.7)
@@
$J\left[\frac{\alpha{p^{2}}\Phi^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)\right]+(1-\alpha)\boldsymbol{x}^{2,(t)}}{\boldsymbol{x}^{k=2,(t)}}\right]=\alpha{p^{2}}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)\right]\cdot D\left[\boldsymbol{\pi}^{2,(t+1)}\right]\cdot J\left[\frac{\boldsymbol{\pi}^{2,(t+1)}}{\boldsymbol{x}^{k,(t)}}\right]+(1-\alpha)I,$ (G.8)
where $J\left[\frac{\boldsymbol{\pi}^{2,(t+1)}}{\boldsymbol{x}^{k,(t)}}\right]$ is given by (G).
@@
Appendix H Proof of Proposition 7
@@
$\left|{\lambda I-J\Phi}\right|=\begin{vmatrix}&\lambda I-X-(1-\alpha)I,&-X&-X\\ &-Y&\lambda I-Y-(1-\alpha)I&-Y\\ &-Z&-Z&\lambda I-Z-(1-\alpha)I\\ \end{vmatrix}=0,$ (H.1)
@@
$\left(\lambda_{i,1}-1+\alpha\right)\left(\lambda_{i,2}-g(\alpha,\hat{\alpha},p^{0},p^{1})\right)=0,$ (H.3)

### v2

@@ Appendix C Derivation of Fact 1
Denote $\boldsymbol{z}^{k,(t)}$ as $\boldsymbol{x}^{k,(t)}-\gamma\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right),k=0,1,2$. For the 0-step travelers, take the Jacobian, $J[\cdot]$, for RHS of Eq. (9) w.r.t. $\boldsymbol{x}^{0,(t)}$:
@@
where $J\left[\frac{\boldsymbol{c}\left(\boldsymbol{\pi}^{1,(t+1)}\right)}{\boldsymbol{x}^{k,(t)}}\right],k=0,1,2,$ is given by Eq. (C).
@@
where $J\left[\frac{\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)}{\boldsymbol{x}^{k,(t)}}\right],k=0,1,2,$ is given by Eq. (C).
@@ Appendix D Proof of Proposition 3
According to Proposition 2, at the DUE, $\boldsymbol{\pi}^{0,(t+1)}=\boldsymbol{\pi}^{1,(t+1)}=\boldsymbol{\pi}^{2,(t+1)}$ and thus the Jacobians of route cost function $D_{2}=D_{1}=D_{0}=D_{\ast}$, where $D_{\ast}$ is the Jacobian of the route cost function evaluated at the DUE $\tilde{\boldsymbol{x}}^{\ast}$. As per the assumption that the projection operator always yields positive route flow, all the $Q$’s and $\hat{Q}$’s in the Jacobian matrix in Eq. (23) reduce to $\bar{Q}$ (defined in Lemma 2). Moreover, since $\alpha=\hat{\alpha}=1$ and $\hat{\gamma}=\gamma$, each sub-block matrix in Eq. (23) reduces to:
@@
Denote $A$ as $\bar{Q}(I-\gamma D_{\ast})$. Thanks to the idempotence of $\bar{Q}$ (Lemma 2), we have $\bar{Q}A=\bar{Q}\bar{Q}(I-\gamma D_{\ast})=A$. With this identity and after performing matrix algebraic calculations, the Jacobian in Eq. (23) can be further represented as:
@@
Using the identity $\bar{Q}A=A$, one can verify that $(\lambda I-A)\cdot(-(A-\bar{Q})A)=(-(A-\bar{Q})A)\cdot(\lambda I-A)$. Hence, $\lambda I-A$ commutes with $-(A-\bar{Q})A$. Referring to Theorem A.2, we have:
@@
Eq. (D) shows that for $|K|=2$, $JP$ has $2\cdot\sum_{w}{|R_{w}|}$ eigenvalues, half of which are the eigenvalues of $\bar{Q}$ and the other half are the ones of $A^{2}$. For each OD pair $w$, we know from Lemma 2 that $\bar{Q}_{w}$ has $|R_{w}|$ eigenvalues with $|R_{w}|-1$ of them being 1 and the other one being 0. So $JP$ will have $\sum_{w}{(|R_{w}|-1)}$ eigenvalues being 1 and $|W|$ ones being 0. Due to the geometric and algebraic multiplicities with eigenvalue one both being $|R_{w}|-1$, every eigenvalue on the unit circle has an associated Jordan block of order 1, which satisfies the condition (ii) in Theorem A.1. As a result, the stability only depends on the eigenvalues of $A^{2}$.
@@
Row and column operations do not change the determinant. Subtracting the third “block column” of the determinant from the first and second block columns and then adding the first and second “block rows” to the third block row, the determinant in Eq. (D.5) becomes an upper triangular matrix:
@@ Appendix E Proof of Lemma 3
$Q\left[\boldsymbol{z}\right](I-\gamma D\left[\boldsymbol{x}\right])$ is not diagonalizable because $Q\left[\boldsymbol{z}\right](I-\gamma D\left[\boldsymbol{x}\right])$ does not commute with $-\gamma Q\left[\boldsymbol{z}\right]D\left[\boldsymbol{x}\right]$. Still, $Q\left[\boldsymbol{z}\right](I-\gamma D\left[\boldsymbol{x}\right])$ share the same eigenvector with $-\gamma Q\left[\boldsymbol{z}\right]D\left[\boldsymbol{x}\right]$ for nonzero eigenvalues. Suppose $\mu_{i}$ is a nonzero eigenvalue of $Q\left[\boldsymbol{z}\right]D\left[\boldsymbol{x}\right]$ with corresponding eigenvector $\boldsymbol{v}$, i.e., $Q\left[\boldsymbol{z}\right]D\left[\boldsymbol{x}\right]\boldsymbol{v}=\mu_{i}\boldsymbol{v},\mu_{i}\neq 0$. Then we have from Eq. (E) that $(-\gamma\xi_{i}\mu_{i})\boldsymbol{v}=(-\gamma\mu_{i}+\gamma^{2}\mu_{i}^{2})\boldsymbol{v}$ where $\xi_{i}$ is the corresponding eigenvalue for $Q\left[\boldsymbol{z}\right](I-\gamma D\left[\boldsymbol{x}\right])$. Solving it yields $\xi_{i}=1-\gamma\mu_{i},\mu_{i}\neq 0$.
@@ Appendix F Proof of Proposition 4
where $A=\bar{Q}(I-\gamma D_{\ast})$ and $\hat{A}=\bar{Q}(I-\hat{\gamma}D_{\ast})$. Using again the idempotence of $\bar{Q}$, one can verify that $(\lambda I-A)$ and $-(A-\bar{Q})\hat{A}$ commute. Referring to Theorem A.2, we have $|(\lambda I-A\hat{A}+\hat{A}-A)||(\lambda I-\bar{Q})|=0$. $A$ and $\hat{A}$ share the same eigenvector $\boldsymbol{v}$ with $\bar{Q}D_{\ast}$ for the eigenvalue $\lambda$ of $\bar{Q}D_{\ast}$ being nonzero. Assume $\bar{Q}D_{\ast}\boldsymbol{v}=\lambda\boldsymbol{v},\lambda\neq 0$. The matrix $A\hat{A}-\hat{A}+A$ determining the stability multiplied by $\boldsymbol{v}$ becomes:
@@
Appendix G Proof of Proposition 5
We prove it by contradiction for each class.
Class-0. Assume $x_{i}^{0,\dagger}>0$ for all routes $i$. The equilibrium condition for class $k=0$ is:
$\boldsymbol{x}^{0,\dagger}=P_{\Omega_{p^{0}}}{\left[\boldsymbol{x}^{0,\dagger}-\gamma\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)\right]}.$ (G.1)
Since $x_{i}^{0,\dagger}>0,\forall i$, the first-order optimality condition implies that for each OD pair $w$, travel costs are equalized across all routes:
$c_{i}(\tilde{\boldsymbol{x}}^{\dagger})=c_{j}(\tilde{\boldsymbol{x}}^{\dagger}),\forall i,j\in R_{w}.$ (G.2)
Two subcases arise: (i) If the DUE requires zero flow on some routes, no $\tilde{\boldsymbol{x}}^{\dagger}\in\Omega_{1}$ can satisfy these equalities, which results in a contradiction; (ii) If the DUE solution requires strictly positive flows on all routes, the equalities reduces to the DUE condition, contradicting the assumption that $\tilde{\boldsymbol{x}}^{\dagger}$ is not DUE. Thus, class $k=0$ must abandon at least one route.
Class-1. Assume $x_{i}^{1,\dagger}>0$ for all routes $i$. The equilibrium condition for class $k=1$ is:
$\boldsymbol{x}^{1,\dagger}=P_{\Omega_{p^{1}}}{\left[\boldsymbol{x}^{1,\dagger}-\gamma\boldsymbol{c}\left(\boldsymbol{\pi}^{1,\dagger}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)\right)\right]},$ (G.3)
where $\boldsymbol{\pi}^{1,\dagger}=\hat{\alpha}P_{\Omega_{1}}{\left[\tilde{\boldsymbol{x}}^{\dagger}-\hat{\gamma}\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)\right]}+(1-\hat{\alpha})\tilde{\boldsymbol{x}}^{\dagger}$.
If all routes in class $k=1$ have positive flow, then the first-order optimality condition implies that for each OD pair $w$:
$c_{i}\left(\boldsymbol{\pi}^{1,\dagger}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)\right)=c_{j}\left(\boldsymbol{\pi}^{1,\dagger}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)\right),\forall i,j\in R_{w}.$ (G.4)
Similar to the analysis of class $k=0$, we consider two subcases.
  1. (i)
    If the DUE admits zero flow on some routes, then no $\boldsymbol{\pi}^{1,\dagger}\in\Omega_{1}$ can satisfy the equalities, leading to a contradiction.
  2. (ii)
    If the DUE admits strictly positive flow on all routes, then $\boldsymbol{\pi}^{1,\dagger}$ is a DUE flow $\tilde{\boldsymbol{x}}^{\ast}$. This case requires more detailed analysis, as there could be many non-DUE total flow patterns ($\tilde{\boldsymbol{x}}^{\dagger}$) that, after the prediction step Eq. (11), become a DUE ($\boldsymbol{\pi}^{1,\dagger}$). In other words, $\boldsymbol{\pi}^{1,\dagger}$ could be the DUE while $\tilde{\boldsymbol{x}}^{\dagger}$ is not. This occurs because the projection operation in Eq. (11) is surjective but not injective.)
    To derive a contradiction, we substitute $\boldsymbol{\pi}^{0,\dagger}=\tilde{\boldsymbol{x}}^{\dagger}$ and $\boldsymbol{\pi}^{1,\dagger}=\tilde{\boldsymbol{x}}^{\ast}$ into the joint VI characterization of the MPE in Eq. (14) with $|K|=2$:
    $\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)^{T}\left(\boldsymbol{x}^{0}-\boldsymbol{x}^{0,\dagger}\right)+\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\ast}\right)^{T}\left(\boldsymbol{x}^{1}-\boldsymbol{x}^{1,\dagger}\right)\geq 0,\forall\boldsymbol{x}^{0}\in\Omega_{p^{0}},\boldsymbol{x}^{1}\in\Omega_{p^{1}}.$ (G.5)
    Since $\tilde{\boldsymbol{x}}^{\dagger}$ is not a DUE, there exists at least one OD pair $w$ with routes $r,s\in R_{w}$ such that $c_{r}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)>c_{s}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)$. Note that $\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\ast}\right)$ is constant across routes in each OD pair, making the term $\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\ast}\right)^{T}\left(\boldsymbol{x}^{1}-\boldsymbol{x}^{1,\dagger}\right)=0$. Therefore, the cost difference must come from class $k=0$ travelers: they must be using at least two routes in OD pair $w$ with unequal costs, i.e., $x_{r}^{0,\dagger}>0,x_{s}^{0,\dagger}>0$ with $c_{r}(\tilde{\boldsymbol{x}}^{\dagger})>c_{s}(\tilde{\boldsymbol{x}}^{\dagger})$.
    We construct a perturbed flow $\boldsymbol{x}^{0}$ by reallocating a small amount $\epsilon>0$ from route $r$ to route $s$, while maintaining the flow on all the other routes:
    $x_{r}^{0}=x_{r}^{0,\dagger}-\epsilon,x_{s}^{0}=x_{s}^{0,\dagger}+\epsilon.$ (G.6)
    For this perturbed flow pattern (which remains in $\Omega_{p^{0}}$), we obtain:
    $\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)^{T}\left(\boldsymbol{x}^{0}-\boldsymbol{x}^{0,\dagger}\right)=\epsilon\left(c_{s}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)-c_{r}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)\right)<0.$ (G.7)
    This inequality shows there exists a flow pattern $\boldsymbol{x}^{0}$ that violates the joint VI of Eq. (G.5), i.e.,
    $\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\dagger}\right)^{T}\left(\boldsymbol{x}^{0}-\boldsymbol{x}^{0,\dagger}\right)+\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{\ast}\right)^{T}\left(\boldsymbol{x}^{1}-\boldsymbol{x}^{1,\dagger}\right)<0,\exists\boldsymbol{x}^{0}\in\Omega_{p^{0}},\boldsymbol{x}^{1}\in\Omega_{p^{1}}.$ (G.8)
    This contradiction completes the proof for the second subcase.
In summary, we have proven that for each traveler class $k\in\{0,1\}$, there exists at least one route $j(k)$ such that $x_{j(k)}^{k,\dagger}=0$.
Appendix H Proof of Example 1
For a DUE with positive flows on both routes, the equilibrium condition requires:
$\tilde{x}_{1}^{\ast}=\frac{a_{2}d+b_{2}-b_{1}}{a_{1}+a_{2}},\tilde{x}_{2}^{\ast}=\frac{a_{1}d+b_{1}-b_{2}}{a_{1}+a_{2}},-a_{1}d<b_{1}-b_{2}<a_{2}d.$ (H.1)
The total cost at DUE equals:
$TC^{\ast}=c_{1}(\tilde{x}_{1}^{\ast})\cdot\tilde{x}_{1}^{\ast}+c_{2}(\tilde{x}_{2}^{\ast})\cdot\tilde{x}_{2}^{\ast}=\frac{a_{1}a_{2}d^{2}+a_{1}b_{2}d+a_{2}b_{1}d}{a_{1}+a_{2}}.$ (H.2)
At non-DUE MPE, Proposition 5 implies each traveler class abandons one route. We first analyze the case where 0-step travelers exclusively use Route 1 (i.e., $x_{1}^{0,\dagger}=p^{0}d$ and $x_{2}^{0,\dagger}=0$). This indicates that $c_{1}(\tilde{x}_{1}^{\dagger})\leq c_{2}(\tilde{x}_{2}^{\dagger})$. For 1-step travelers, $\boldsymbol{x}^{1,\dagger}$ can only take two possible values: $(0,p^{1}d)^{T}$ or $(p^{1}d,0)^{T}$. In the latter case, $c_{1}(\tilde{x}_{1}^{\dagger})=a_{1}d+b_{1}$ and $c_{2}(\tilde{x}_{2}^{\dagger})=b_{2}$. Due to the third inequality of Eq. (H.1), we have $c_{1}(\tilde{x}_{1}^{\dagger})>c_{2}(\tilde{x}_{2}^{\dagger})$, which contradicts $c_{1}(\tilde{x}_{1}^{\dagger})\leq c_{2}(\tilde{x}_{2}^{\dagger})$. Therefore, $\boldsymbol{x}^{1,\dagger}$ must equal $(0,p^{1}d)^{T}$. As a result, we can deduce that $\boldsymbol{x}^{0,\dagger}=(p^{0}d,0)$ and $\boldsymbol{x}^{1,\dagger}=(0,p^{1}d)$ constitute a non-DUE MPE.
At the non-DUE MPE of $\boldsymbol{x}^{0,\dagger}=(p^{0}d,0)$ and $\boldsymbol{x}^{1,\dagger}=(0,p^{1}d)$, $c_{1}(\tilde{x}_{1}^{\dagger})\leq c_{2}(\tilde{x}_{2}^{\dagger})$ indicates that $a_{1}p^{0}d+b_{1}\leq a_{2}(1-p^{0})d+b_{2}$, which simplifies to $p^{0}\leq\frac{b_{2}-b_{1}+a_{2}d}{(a_{1}+a_{2})d}$. The total cost at this non-DUE MPE becomes:
$TC^{\dagger}=(a_{1}p^{0}d+b_{1})p^{0}d+(a_{2}(1-p^{0})d+b_{2})(1-p^{0})d.$ (H.3)
The difference between $TC^{\dagger}$ and $TC^{\ast}$ can be expressed as a quadratic function of $p^{0}$:
$\Delta TC=TC^{\dagger}-TC^{\ast}=(a_{1}+a_{2})d^{2}\cdot(p^{0})^{2}-[2a_{2}d^{2}-(b_{1}-b_{2})d]\cdot p^{0}+\frac{a_{2}d(a_{2}d+b_{2}-b_{1})}{a_{1}+a_{2}}.$ (H.4)
$\Delta TC$ has two roots: $\frac{a_{2}}{a_{1}+a_{2}}$ and $\frac{b_{2}-b_{1}+a_{2}d}{(a_{1}+a_{2})d}$. Since the coefficient $(a_{1}+a_{2})d^{2}$ is positive ($a_{1},a_{2}>0$), the parabola opens upward. If $b_{2}<b_{1}$, the left root is $\frac{b_{2}-b_{1}+a_{2}d}{(a_{1}+a_{2})d}$. Any $p^{0}\leq\frac{b_{2}-b_{1}+a_{2}d}{(a_{1}+a_{2})d}$ would lie to the left of this root, resulting in $TC^{\dagger}>TC^{\ast}$. If $b_{2}>b_{1}$, the right root is $\frac{b_{2}-b_{1}+a_{2}d}{(a_{1}+a_{2})d}$. Any value $p^{0}\in\left(\frac{a_{2}}{a_{1}+a_{2}},\frac{b_{2}-b_{1}+a_{2}d}{(a_{1}+a_{2})d}\right)$ would yield a lower $TC^{\dagger}$ than $TC^{\ast}$.
The derivation for the case where 0-step travelers use only route 2 follows analogous logic. In this case, $\boldsymbol{x}^{0,\dagger}=(0,p^{0}d)$ and $\boldsymbol{x}^{1,\dagger}=(p^{1}d,0)$ constitute the other non-DUE MPE. Through similar deduction logic, we can prove that at this second non-DUE MPE, $TC^{\dagger}<TC^{\ast}$ when $b_{2}<b_{1}$ and $p^{1}\in\left(\frac{b_{2}-b_{1}+a_{2}d}{(a_{1}+a_{2})d},\frac{a_{2}}{a_{1}+a_{2}}\right)$.
Appendix I Derivation of Fact 2
@@
For the 0-step travelers, take the Jacobian, $J[\cdot]$, of the dynamical system in Eq. (26) w.r.t. $\boldsymbol{x}^{k,(t)}$:
@@
$=\alpha p^{0}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\alpha)I,$ (I.1)
@@
$=\alpha p^{0}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D[\tilde{\boldsymbol{x}}^{(t)}].$ (I.2)
@@
$=D\left[\boldsymbol{\pi}^{1,(t+1)}\right]\cdot\left(\hat{\alpha}\Upsilon^{\hat{\theta}}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\hat{\alpha})I\right).$ (I.3)
@@
$\quad=\alpha p^{1}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{1,(t+1)}\right)\right]\cdot D\left[\boldsymbol{\pi}^{1,(t+1)}\right]\cdot\left(\hat{\alpha}\Upsilon^{\hat{\theta}}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\hat{\alpha})I\right)+(1-\alpha)I,$ (I.4)
@@
$\quad=\alpha{p^{1}}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{1,(t+1)}\right)\right]\cdot D\left[\boldsymbol{\pi}^{1,(t+1)}\right]\cdot\left(\hat{\alpha}\Upsilon^{\hat{\theta}}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]\cdot D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\hat{\alpha})I\right).$ (I.5)
@@
$\cdot\left(D\left[\boldsymbol{\pi}^{1,(t+1)}\right]\cdot\left(\hat{\alpha}\Upsilon^{\hat{\theta}}\left[\boldsymbol{c}\left(\tilde{\boldsymbol{x}}^{(t)}\right)\right]D\left[\tilde{\boldsymbol{x}}^{(t)}\right]+(1-\hat{\alpha})I\right)\right)+(1-\hat{\alpha})I.$ (I.6)
@@
$J\left[\frac{\alpha{p^{2}}\Phi^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)\right]+(1-\alpha)\boldsymbol{x}^{2,(t)}}{\boldsymbol{x}^{k\neq 2,(t)}}\right]=\alpha{p^{2}}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)\right]\cdot D\left[\boldsymbol{\pi}^{2,(t+1)}\right]\cdot J\left[\frac{\boldsymbol{\pi}^{2,(t+1)}}{\boldsymbol{x}^{k,(t)}}\right],$ (I.7)
@@
$J\left[\frac{\alpha{p^{2}}\Phi^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)\right]+(1-\alpha)\boldsymbol{x}^{2,(t)}}{\boldsymbol{x}^{k=2,(t)}}\right]=\alpha{p^{2}}\Upsilon^{\theta}\left[\boldsymbol{c}\left(\boldsymbol{\pi}^{2,(t+1)}\right)\right]\cdot D\left[\boldsymbol{\pi}^{2,(t+1)}\right]\cdot J\left[\frac{\boldsymbol{\pi}^{2,(t+1)}}{\boldsymbol{x}^{k,(t)}}\right]+(1-\alpha)I,$ (I.8)
where $J\left[\frac{\boldsymbol{\pi}^{2,(t+1)}}{\boldsymbol{x}^{k,(t)}}\right]$ is given by Eq. (I).
@@
Appendix J Proof of Proposition 8
@@
$\left|{\lambda I-J\Phi}\right|=\begin{vmatrix}&\lambda I-X-(1-\alpha)I,&-X&-X\\ &-Y&\lambda I-Y-(1-\alpha)I&-Y\\ &-Z&-Z&\lambda I-Z-(1-\alpha)I\\ \end{vmatrix}=0,$ (J.1)
@@
$\left(\lambda_{i,1}-1+\alpha\right)\left(\lambda_{i,2}-g(\alpha,\hat{\alpha},p^{0},p^{1})\right)=0,$ (J.3)